# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

### v1

@@ Abstract.
2020 Mathematics Subject Classification: Primary 30D35; Secondary 32Q45 and 32H30
The research of Min Ru is supported in part by Simon Foundations grant award #531604 and #521604. The second-named author was supported in part by Taiwan’s NSTC grant 113-2115-M-001-011-MY3.
@@ 1. Introduction
Let $X$ be an algebraic variety, and $D$ an effective divisor (possibly empty). The set $X\backslash D$ is Brody hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is constant. It is Brody quasi-hyperbolic if every entire curve $f:\mathbb{C}\to X\backslash D$ is degenerate, meaning its image is contained a proper Zariski-closed subset $Z$ of $X$. We call such $Z$ the exceptional set. If $Z$ is independent of $f$, we call $X\backslash D$ Brody strong quasi-hyperbolic. This leads us to the Griffiths-Lang’s conjecture.
Griffiths-Lang’s conjecture. Let $X$ be a smooth complex projective variety, $D$ a normal crossing divisor on $X$, and $K_{X}$ the canonical divisor on $X$. If $K_{X}+D$ is big, then $X\backslash D$ is Brody quasi-hyperbolic.
@@
holds for any $\epsilon>0$, where notation $\leq_{\rm exc}$ means the inequality holds for all $r\in(0,\infty)$ except for a set of finite measure.
@@
$N_{f}(D,r)$ is the counting function with respect to $D$ given by $N_{f}^{(\infty)}(D,r)$, and $T_{D,f}(r)$ is the (Nevanlinna) height function relative to the divisor $D$ (referring to [25, Section 12]).
Paul Vojta formulated the following general abc conjecture (see [25, Conjecture 15.2] and [25, Conjecture 23.4]), which extends the above Griffiths’ conjecture. It is also viewed as the quantitive version of Campana’s orbifold conjecture which will be stated later in the introduction.
Vojta’s abc conejcture. For any $\epsilon>0$, there exists a proper Zariski-closed subset $Z$ of $X$ such that for any entire curve $f:\mathbb{C}\to X$ whose image is not contained in $Z$, the inequality
@@
holds, where $A$ is an ample divisor on $X$.
Note that Vojta’s abc conjecture strengthens Griffiths’ conjecture in two key ways: (i) the exceptional set is independent of the map $f$, and (ii) the counting function is truncated at level one. Proving these two features is quite challenging, even if the aforementioned version of Griffiths’ conjecture could be established. In this paper, we first prove Vojta’s abc conjecture for entire curves in ${\mathbb{P}}^{n}({\mathbb{C}})$, assuming the map is highly ramified over the coordinate hyperplanes. The precise statement of our main result is as follows. Additionally, we highlight that the exceptional set obtained in the theorem can be determined explicitly-this is a challenging issue in the study of Nevanlinna theory and complex hyperbolicity.
@@
Note that the case when $n=2$ is due to Ji Guo and the second named author ([9, Theorem 1.4]). The new idea in this paper is motivated by the recent paper [2] by considering the parabolic Riemann surface $\mathcal{Y}_{f}:=\mathbb{C}\setminus{f^{-1}(D)}$, reducing the problem to the omitting case. To achieve this, we need to extend the results in [9] and [7] to the parabolic setting.
@@
    The Campana’s conjecture is a consequence of the above Vojta’s strong abc conecutre.
@@
    Let $D:={\rm Supp}(\Delta)=D_{1}+\cdots+D_{q}$. Recall that the pair $(X,D)$ is said to be of log-general type if the divisor $K_{X}+D$ is big. It’s clear that $(X,D)$ is of log-general type if $(X,\Delta)$ is of general type, i.e. $K_{X}+\Delta$ is big. On the other hand, the condition that $(X,D)$ is of log-general type implies that $(X,\Delta)$ is of general type if $(X,\Delta)$ has a sufficiently large multiplicity along each $D_{i}$, $1\leq i\leq q$. ((((See [12, Corollary 2.2.24].))))
@@
Let $\Delta_{0}$ be an orbifold divisor of ${\mathbb{P}}^{n}({\mathbb{C}})$ and let $H_{0},H_{1},\dots,H_{n}$ be the coordinate hyperplanes in ${\mathbb{P}}^{n}({\mathbb{C}})$, such that $\Delta_{0}$ and $H_{0},H_{1},\dots,H_{n}$ are in general position. Let $m_{i}\in(1,\infty)\cap{\mathbb{Q}}$, $0\leq i\leq n$, and $\Delta=\Delta_{0}+(1-{1\over m_{0}})H_{0}+(1-{1\over m_{1}})H_{1}+\cdots+(1-{1\over m_{n}})H_{n}$. Assume that $\deg\Delta>n+1$. Then there exist a proper Zariski closed subset $W$ of ${\mathbb{P}}^{n}({\mathbb{C}})$ and an effectively computable positive integer $\ell$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{0},m_{1},\dots,m_{n}\}\geq\ell$ must be contained in $W$.
The next case involves $n+1$ components, not all lines, in $\mathbb{P}^{n}({\mathbb{C}})$.
@@
Let $F_{i}$, $11\leq i\leq n+$ 1, be homogeneous irreducible polynomials of positive degrees in $\mathbb{C}[x_{0},\ldots,x_{n}]$. Assume that $D_{i}:=[F_{i}=0]\subset\mathbb{P}^{n}(\mathbb{C})$, $11\leq i\leq n+$ 1, intersect transversally. Let $\Delta=(1-{1\over m_{1}})D_{1}+\cdots+(1-{1\over m_{n+1}})D_{n+1}$ be an orbifold divisor of $\mathbb{P}^{n}({\mathbb{C}})$. Assume that $\deg\Delta>n+1$. Then, there exist an effectively computable positive integer $\ell$ and a proper Zariski closed subset $W$ of $\mathbb{P}^{n}({\mathbb{C}})$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{1},\dots,m_{n+1}\}\geq\ell$ must be contained in $W$.
Note that $\mathbb{P}^{n}({\mathbb{C}})$ is a toric variety when we identify ${\mathbb{G}}_{m}^{n}$ with ${\mathbb{P}}^{n}({\mathbb{C}})\backslash\cup_{i=0}^{n}H_{i}$, where $H_{0},\dots,H_{n}$ are the $n+1$ coordinate hyperplanes. A toric variety $X$ is defined as a variety over ${\mathbb{C}}$ that contains ${\mathbb{G}}_{m}^{n}$ as a dense open subvariety, with the action of ${\mathbb{G}}_{m}^{n}$ on itself extends to an algebraic action of ${\mathbb{G}}_{m}^{n}$ on $X$. We refer to [5] for more details on toric varieties. Using the arguments in [13] and [6], we can extend our Main Theorem to projective toric varieties. Recall from [13] that an “admissible pair” is a couple $(X,V)$ where $V$ is a nonsingular variety embedded in a nonsingular projective variety $X$ in such a way that $D_{0}=X\setminus V$ is a normal crossings divisor. A Campana orbifold $(X,\Delta)$ is said to be associated to the admissible pair $(X,V)$ if $D_{0}$ is the support of $\Delta$.
@@
Let $X$ be a nonsingular projective toric variety and let $D$ be an effective reduced divisor on the admissible pair $(X,\mathbb{G}_{m}^{n})$, such that $D$ is in general position with the boundary $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $A$ be a big divisor on $X$. For every $\epsilon>0$ there is a positive integral number $\ell$ and a proper Zariski-closed subset $Z$ of $X$ such that, for any Campana orbifold $(X,\Delta)$ associated to the admissible pair $(X,\mathbb{G}_{m}^{n})$ with multiplicities at least $\ell$ along the boundary $D_{0}$, the following two inequalities hold:
@@
Let $X$ be a nonsingular projective toric variety, and let $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $Y$ be a nonsingular complex projective variety with a finite morphism $\pi:Y\to X$. Let $H:=\pi^{*}(D_{0})$ and $R\subset Y$ the ramification divisor of $\pi$ omitting components from the support of $H$. Assume that $\pi(R)$ and $D_{0}$ are in general position on $X$. Then there exists a positive integer $\ell$ such that if the orbifold pair $(Y,\Delta)$ is of general type with ${\rm Supp}(\Delta)={\rm Supp}(H)$ and multiplicity at least $\ell$ for each component, then there exists a proper Zariski closed subset $W$ of $Y$ such that the image of any non-constant orbifold entire curve $f:{\mathbb{C}}\rightarrow(Y,\Delta)$ must be contained in $W$.
In Section 2, we prove a version of the Main Theorem for rational maps and explain the motivation for using parabolic Riemann surfaces. Section 3 reviews basic results from Nevanlinna theory relevant to parabolic settings, as well as the key theorems from [2]. Furthermore, we extend the GCD theorem developed in [14] and in [9] to parabolic setting. In Section 4, we establish the abc theorem for parabolic Riemann surfaces, building on the methods developed in [7]. Section 5 presents the proof of the Main Theorem, while Section 6 extends results to toric varieties.
@@ 2. Motivation through the rational case
In [7], a version of Vojta’s abc conjecture for algebraic tori over algebraic function fields is established. As we will see, the algebraic reduction method presented in this article is a key technical component in proving the Main Theorem. The core idea behind the proof is also inspired by the function field approach. In this section, we will prove a version of the Main Theorem for rational functions and explain the role of parabolic Riemann surfaces in this context.
@@ 2.1. The Rational Case
The statements and arguments apply verbatim to algebraic function fields. Here, let $K=\mathbb{C}(t)$ be the function field of $\mathbb{P}^{1}({\mathbb{C}})$. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. We denote by ${\mathcal{O}}_{S}^{*}$ the set of $S$-units, which consists of rational functions with no zeros and poles outside of $S$. For a rational function $f=\frac{P}{Q}$, where $P$ and $Q$ are coprime polynomials, we define $\deg f:=\max\{\deg P,\deg Q\}$ and denote by ${\rm ord}_{\mathbf{p}}^{+}(f)$ the order of the zero of $f$ at $\mathbf{p}\in\mathbb{P}^{1}(\mathbb{C})$. We use the notation $N_{S}(f)$ (respectively $N_{S}^{(1)}(f)$) to count the number of zeros of $f$ with multiplicity (respectively, without multiplicity) outside of $S$. When $S$ is empty, we simply write $N(f)$ (respectively $N^{(1)}(f)$) to count the number of zeros with multiplicity (respectively, without multiplicity) on $\mathbb{P}^{1}(\mathbb{C})$.
@@
Here $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ and can be determined explicitly and (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
@@
for any non-constant entire rational curve $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, provided that $\mathbf{f}(\mathbb{C})\not\subset Z$ and ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$, and for all $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$. Here, $\deg\mathbf{f}=\max\{\deg f_{0},\ldots,\deg f_{n}\}$.
@@
Denote by $\ell$ a sufficiently large integer to be determined later. Let $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros such that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$. Let
@@
and $f_{i}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$ for $0\leq i\leq n$. Let $\epsilon>0$ be given. We now apply Theorem 5 to find a positive integer $c_{0}$ and a proper Zariski closed subset $Z_{0}$ independent of $\tilde{S}_{\mathbf{f}}$ of $\mathbb{A}^{n}$ such that we have either the inequality
@@
(2.4) $N_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\leq\epsilon\deg\mathbf{f},$
(2.5) $N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f},$
if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$, where $\tilde{Z}_{0}$ is the projective closure of $Z_{0}$ in $\mathbb{P}^{n}({\mathbb{C}})$. Here $G_{0}(x_{1},\ldots,x_{n}):=G(1,x_{1},\ldots,x_{n})$ and $u_{i}=\frac{f_{i}}{f_{0}}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$, $1\leq i\leq n$. Since the constant $c_{0}$ and the exceptional set $Z$ in Theorem 5 do not depend on $S$, it is important to note that $c_{0}$ and $Z_{0}$ do not depend on $\mathbf{f}$ as well. We first show that (2.3) does not hold if $\ell>2c_{0}n$. Indeed, if (2.3) holds with $\ell>2c_{0}n$, then it follows from (2.2) and the choice of $\ell$ that (2.3) yields
@@
Therefore, $\max_{0\leq i\leq n}\{\deg f_{i}\}<2c_{0}.$ This is not possible since not all the polynomials $f_{i}$ are constant, the assumption that ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$ if ${\rm ord}_{\mathbf{p}}(f_{i})>0$ for $0\leq i\leq n$ and $\mathbf{p}\in{\mathbb{P}}^{1}(\mathbb{C})$ implies that $2nc_{0}<\ell<\max_{0\leq i\leq n}\{\deg f_{i}\}$.
We continue to let $\ell>2c_{0}n$. As we have shown that (2.3) does not hold, it follows that (2.4) and (2.5) both hold if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$. Since $N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G_{0}(u_{1},\ldots,u_{n}))$, it follows trivially from (2.5) that
(2.6) $N^{(1)}(G(\mathbf{f}))\geq N^{(1)}(G_{0}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f}.$
@@
It remains to show (i). We now compute zeros for points in $\tilde{S}_{\mathbf{f}}$. We will use Propostion 7. Let $W$ be the Zariski closed subset of $\mathbb{P}^{n}(\mathbb{C})$ containing all hypersurfaces defined by all possible subsums (including $G$) in the expansion of $G$. Assume that $\mathbf{f}(\mathbb{C})\not\subset W$. Since $G_{0}(0,\dots,0)=G(1,0,\ldots,0)\neq 0$, we now apply Propositon 7 to get
$\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\},$
@@
$m_{\tilde{S}_{\mathbf{f}}}(G_{0}(u_{1},\ldots,u_{n})):=\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(G_{0}(u_{1},\ldots,u_{n}))\leq\frac{(n+1)\tilde{c}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\epsilon\deg\mathbf{f}$

### v2

@@ Abstract.
2020 Mathematics Subject Classification: Primary 32H30; Secondary 32Q45 and 30D35
The research of Min Ru is supported in part by Simon Foundations grant award #531604 and #521604. The research of Julie Tzu-Yueh Wang was supported in part by Taiwan’s NSTC grant 113-2115-M-001-011-MY3.
@@ 1. Introduction
[…]
@@
holds for any $\epsilon>0$, where $A$ is an ample divisor on $X$, and the notation $\leq_{\rm exc}$ means the inequality holds for all $r\in(0,\infty)$ except for a set of finite measure.
@@
$N_{f}(D,r)$ is the counting function with respect to $D$ given by $N_{f}^{(\infty)}(D,r)$, and $T_{D,f}(r)$ is the (Nevanlinna) height function relative to the divisor $D$ (referring to [27, Section 12]).
Paul Vojta formulated the following general abc conjecture (see [27, Conjecture 15.2] and [27, Conjecture 23.4]), which extends the above Griffiths’ conjecture. It is also viewed as the quantitive version of Campana’s orbifold conjecture which will be stated later in the introduction.
Vojta’s abc conjecture. Let $X$ be a smooth complex projective variety, $D$ a normal crossing divisor on $X$, and $K_{X}$ a canonical divisor on $X$. For any $\epsilon>0$, there exists a proper Zariski-closed subset $Z$ of $X$ such that for any entire curve $f:\mathbb{C}\to X$ whose image is not contained in $Z$, the inequality
@@
holds for any $\epsilon>0$, where $A$ is an ample divisor on $X$.
Note that Vojta’s abc conjecture strengthens Griffiths’ conjecture in two key ways: (i) the exceptional set $Z$ above is independent of the map $f$, and (ii) the counting function is truncated at level one. Proving these two features is quite challenging, even if the aforementioned version of Griffiths’ conjecture could be established. In this paper, we first prove Vojta’s abc conjecture for entire curves in ${\mathbb{P}}^{n}({\mathbb{C}})$, assuming the map is highly ramified over the coordinate hyperplanes. The precise statement of our main result is stated below. Additionally, we highlight that the exceptional set obtained in the theorem can be determined explicitly-this is a challenging issue in the study of Nevanlinna theory and complex hyperbolicity.
@@
We note that this theorem corresponds to a case of Vojta’s abc conjecture for entire curves with $D=H_{0}+\cdots+H_{n}+[G=0]$, so that $K_{\mathbb{P}^{n}+D}$ is linearly equivalent to $[G=0]$. The case $n=2$ was previously established by Ji Guo and the second named author ([9, Theorem 1.4]). The new idea in this paper is motivated by the recent paper [2], where the authors considered the parabolic Riemann surface $\mathcal{Y}_{f}:=\mathbb{C}\setminus{f^{-1}(D)}$, reducing the problem to the omitting case. To achieve this in our setting, we need to extend the results in [7] and [9] to the parabolic setting.
@@
    The Campana’s conjecture is a consequence of the above Vojta’s abc conjecutre.
@@
    Given an orbifold divisor $\Delta$ on $X$, let $D:={\rm Supp}(\Delta)=D_{1}+\cdots+D_{q}$. Recall that the pair $(X,D)$ is said to be of log-general type if the divisor $K_{X}+D$ is big. It is clear that $(X,D)$ is of log-general type if $(X,\Delta)$ is of general type, i.e. $K_{X}+\Delta$ is big. On the other hand, the condition that $(X,D)$ is of log-general type implies that $(X,\Delta)$ is of general type if $(X,\Delta)$ has a sufficiently large multiplicity along each $D_{i}$, $1\leq i\leq q$. ((See [11, Corollary 2.2.24].))
@@
Let $\Delta_{0}$ be an orbifold divisor on ${\mathbb{P}}^{n}({\mathbb{C}})$ and let $H_{0},H_{1},\dots,H_{n}$ be the coordinate hyperplanes in ${\mathbb{P}}^{n}({\mathbb{C}})$, such that $\Delta_{0}$ and $H_{0},H_{1},\dots,H_{n}$ are in general position. Let $m_{i}\in(1,\infty)\cap{\mathbb{Q}}$, $0\leq i\leq n$, and $\Delta=\Delta_{0}+(1-{1\over m_{0}})H_{0}+(1-{1\over m_{1}})H_{1}+\cdots+(1-{1\over m_{n}})H_{n}$. Assume that $\deg\Delta>n+1$. Then there exist a proper Zariski closed subset $W$ of ${\mathbb{P}}^{n}({\mathbb{C}})$ and an effectively computable positive integer $\ell$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{0},m_{1},\dots,m_{n}\}\geq\ell$ must be contained in $W$.
The next case involves $n+1$ components, not all are hyperplanes, in $\mathbb{P}^{n}({\mathbb{C}})$.
@@
Let $F_{i}$, $1\leq i\leq n+1$, be homogeneous irreducible polynomials of positive degrees in $\mathbb{C}[x_{0},\ldots,x_{n}]$. Assume that $D_{i}:=[F_{i}=0]\subset\mathbb{P}^{n}(\mathbb{C})$, $1\leq i\leq n+1$, intersect transversally. Let $\Delta=(1-{1\over m_{1}})D_{1}+\cdots+(1-{1\over m_{n+1}})D_{n+1}$ be an orbifold divisor of $\mathbb{P}^{n}({\mathbb{C}})$. Assume that $\deg\Delta>n+1$. Then, there exist an effectively computable positive integer $\ell$ and a proper Zariski closed subset $W$ of $\mathbb{P}^{n}({\mathbb{C}})$ such that the image of any orbifold entire curve $f:{\mathbb{C}}\rightarrow({\mathbb{P}}^{n}({\mathbb{C}}),\Delta)$ with $\mathop{\mathrm{min}}\{m_{1},\dots,m_{n+1}\}\geq\ell$ must be contained in $W$.
[…]
@@
Let $X$ be a nonsingular projective toric variety and let $D$ be an effective reduced divisor on the admissible pair $(X,\mathbb{G}_{m}^{n})$, such that $D$ is in general position with the boundary $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $A$ be a big divisor on $X$. Then, for every $\epsilon>0$, there is a positive integer $\ell$ and a proper Zariski-closed subset $Z$ of $X$ such that, for any Campana orbifold $(X,\Delta)$ associated to the admissible pair $(X,\mathbb{G}_{m}^{n})$ with multiplicities at least $\ell$ along the boundary $D_{0}$, the following two inequalities hold:
@@
Let $X$ be a nonsingular projective toric variety, and let $D_{0}=X\setminus\mathbb{G}_{m}^{n}$. Let $Y$ be a nonsingular complex projective variety with a finite morphism $\pi:Y\to X$. Let $H:=\pi^{*}(D_{0})$ and $R\subset Y$ the ramification divisor of $\pi$ omitting components from the support of $H$. Assume that $\pi(R)$ and $D_{0}$ are in general position on $X$. Then there exists a positive integer $\ell$ such that, if the orbifold pair $(Y,\Delta)$ is of general type with ${\rm Supp}(\Delta)={\rm Supp}(H)$ and multiplicity at least $\ell$ for each component, then there exists a proper Zariski closed subset $W$ of $Y$ such that the image of any non-constant orbifold entire curve $f:{\mathbb{C}}\rightarrow(Y,\Delta)$ must be contained in $W$.
The paper is organized as follows. In Section 2, we prove a version of the Main Theorem for rational maps and explain the motivation for using parabolic Riemann surfaces. Section 3 reviews basic results from Nevanlinna theory relevant to parabolic settings, as well as the key theorems from [2]. Furthermore, we extend the GCD theorem developed in [13] and in [9] to parabolic setting. In Section 4, we establish the abc theorem for parabolic Riemann surfaces, building on the methods developed in [7]. Section 5 presents the proof of the Main Theorem, while Section 6 extends results to toric varieties.
@@ 2. Motivation through the rational case
In [7], a version of Vojta’s abc conjecture for algebraic tori over algebraic function fields is established. As we will see, the algebraic reduction method presented in [7] is a key technical component in proving our Main Theorem. The core idea behind the proof is also inspired by the function field approach. In this section, we will prove a version of the Main Theorem for rational functions and explain the role of parabolic Riemann surfaces in this context.
@@ 2.1. The Rational Case
The statements and arguments apply verbatim to algebraic function fields (see also [6, Theorem 1.13]). Here, let $K=\mathbb{C}(t)$ be the function field of $\mathbb{P}^{1}({\mathbb{C}})$. Let $S$ be a finite subset of $\mathbb{P}^{1}(\mathbb{C})$. We denote by ${\mathcal{O}}_{S}^{*}$ the set of $S$-units, which consists of rational functions with no zeros and poles outside of $S$. For a rational function $f=\frac{P}{Q}$, where $P$ and $Q$ are coprime polynomials, we define $\deg f:=\max\{\deg P,\deg Q\}$ and denote by ${\rm ord}_{\mathbf{p}}^{+}(f)$ the order of the zero of $f$ at $\mathbf{p}\in\mathbb{P}^{1}(\mathbb{C})$. We use the notation $N_{S}(f)$ (respectively $N_{S}^{(1)}(f)$) to count the number of zeros of $f$ with multiplicity (respectively, without multiplicity) outside of $S$. When $S$ is empty, we simply write $N(f)$, which equals $\deg f$, to count the number of zeros with multiplicity on $\mathbb{P}^{1}(\mathbb{C})$; $N^{(1)}(f)$ denotes the same count without multiplicity.
@@
Here $c_{0}$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$. Moreover, the exceptional set $Z$ can be expressed as the zero locus of a finite set $\Sigma\subset\mathbb{C}[x_{1},\ldots,x_{n}]$ with the following properties: (Z1) $\Sigma$ depends on $\epsilon$ and $G$ but independent of $S$ and can be determined explicitly and (Z2) $|\Sigma|$ and the degree of each polynomial in $\Sigma$ can be effectively bounded from above in terms of $\epsilon$, $n$, and the degree of $G$.
The construction of the set of polynomials $\Sigma$ follows from the specialization lemmas [7, Lemma 15 and 16]. In particular, it is evident that $\Sigma$ is independent of the choice of $S$. Note that $\mathbf{u}:=(u_{1},\ldots,u_{n})\in({\mathcal{O}}_{S}^{*})^{n}\setminus Z$ means that $F(\mathbf{u})\neq 0$ (a nontrivial function and hence not identically zero) for every $F\in\Sigma$. This formulation is closely analogous to the number field case. Alternatively, one may view $\mathbf{u}$ as a morphism: $\mathbb{P}^{1}(\mathbb{C})\setminus S\to\mathbb{C}^{n}$; then the condition $F(\mathbf{u})\neq 0$ for every $F\in\Sigma$ is equivalent to saying that the image $\mathbf{u}(\mathbb{P}^{1}(\mathbb{C})\setminus S)$ is not contained in $Z$.
@@
for any non-constant entire rational curve $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, provided that $\mathbf{f}(\mathbb{C})\not\subset Z$ and that the following condition is satisfied: for each $0\leq i\leq n$ and $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$, if ${\rm ord}_{\mathbf{p}}(f_{i})>0$, then ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$. Here, $\deg\mathbf{f}=\max\{\deg f_{0},\ldots,\deg f_{n}\}$.
@@
The proof of Theorem 6 (i) can be outline as the follows. Given a rational map $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros, we define a set $S_{\mathbf{f}}$ consisting of the zeros of all $f_{i}$. With this choice, each coordinate function $f_{i}$ becomes an $S_{\mathbf{f}}$-unit, so that Theorem 5 (ii) can be applied to estimate the multiplicities of the zeros of $G(\mathbf{f})$ outside $S_{\mathbf{f}}$. We then use Proposition 7 to estimate the zeros of $G(\mathbf{f})$ inside $S_{\mathbf{f}}$. Crucially, the constants and exceptional sets appearing in Theorem 5 and Proposition 7 do not depend on $S$, allowing this method to be used for all rational maps.
@@
Denote by $\ell$ a sufficiently large integer to be determined later. Let $\mathbf{f}=(f_{0},\ldots,f_{n})$, where $f_{0},\ldots,f_{n}\in\mathbb{C}[z]$ are polynomials without common zeros. By the assumption, for each $0\leq i\leq n$ and $\mathbf{p}\in\mathbb{P}^{1}({\mathbb{C}})$, if ${\rm ord}_{\mathbf{p}}(f_{i})>0$, then ${\rm ord}_{\mathbf{p}}(f_{i})\geq\ell$. Let
@@
and $f_{i}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$ for $0\leq i\leq n$. Let $\widetilde{G}(x_{1},\ldots,x_{n}):=G(1,x_{1},\ldots,x_{n})$ and $u_{i}=\frac{f_{i}}{f_{0}}\in\mathcal{O}_{\tilde{S}_{\mathbf{f}}}^{*}$, $1\leq i\leq n$. Then $\widetilde{G}(u_{1},\ldots,u_{n}):=f_{0}^{-\deg G}G({\mathbf{f}}).$ An elementary argument shows that
(2.3) $\max_{1\leq i\leq n}\{\deg u_{i}\}\leq\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\sum_{j=1}^{n}\deg u_{j}.$
Let $\epsilon>0$ be given. We now apply Theorem 5 for $\widetilde{G}\in\mathbb{C}[x_{1},\ldots,x_{n}]$ and $\tilde{S}_{\mathbf{f}}$ to find a positive integer $c_{0}$ and a proper Zariski closed subset $Z_{0}$ of $\mathbb{A}^{n}$ independent of $\tilde{S}_{\mathbf{f}}$ such that we have either the inequality
@@
(2.5) $N_{\tilde{S}_{\mathbf{f}}}(G({\mathbf{f}}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G({\mathbf{f}}))=N_{\tilde{S}_{\mathbf{f}}}(\widetilde{G}(u_{1},\ldots,u_{n}))-N_{\tilde{S}_{\mathbf{f}}}^{(1)}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq\epsilon\deg\mathbf{f},$
(2.6) $N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G({\mathbf{f}}))=N_{\tilde{S}_{\mathbf{f}}}^{(1)}(\widetilde{G}(u_{1},\ldots,u_{n}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f},$
if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$, where $\tilde{Z}_{0}$ is the projective closure of $Z_{0}$ in $\mathbb{P}^{n}({\mathbb{C}})$. Since the constant $c_{0}$ and the exceptional set $Z$ in Theorem 5 do not depend on $S$, it is important to note that $c_{0}$ and $\tilde{Z}_{0}$ do not depend on $\mathbf{f}$ as well. We first show that (2.4) does not hold if $\ell>2c_{0}n$. Indeed, if (2.4) holds with $\ell>2c_{0}n$, then it follows from (2.2) and the choice of $\ell$ that (2.4) yields
@@
Therefore, $\max_{0\leq i\leq n}\{\deg f_{i}\}<2c_{0}.$ This is not possible since the $f_{i}$ are not all constant, and by our assumption that every zero has multiplicity at least $\ell$, we have $2nc_{0}<\ell<\max_{0\leq i\leq n}\{\deg f_{i}\}$.
We continue to let $\ell>2c_{0}n$. As we have shown that (2.4) does not hold, it follows that (2.5) and (2.6) both hold if $\mathbf{f}(\mathbb{C})\not\subset\tilde{Z}_{0}$. Then it follows trivially from (2.6) that
(2.7) $N^{(1)}(G(\mathbf{f}))\geq N_{\tilde{S}_{\mathbf{f}}}^{(1)}(G({\mathbf{f}}))\geq(\deg G-\epsilon)\cdot\deg\mathbf{f}.$
@@
It remains to show (i). We now compute zeros for points in $\tilde{S}_{\mathbf{f}}$. We will use Propostion 7. Let $W$ be the Zariski closed subset of $\mathbb{P}^{n}(\mathbb{C})$ containing all hypersurfaces defined by all possible subsums (including $G$) in the expansion of $G$. Assume that $\mathbf{f}(\mathbb{C})\not\subset W$. Since $\widetilde{G}(0,\dots,0)=G(1,0,\ldots,0)\neq 0$, we now apply Propositon 7 to get
$\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq\tilde{c}\max\{1,|\tilde{S}_{\mathbf{f}}|-2\},$
@@
$m_{\tilde{S}_{\mathbf{f}}}(\widetilde{G}(u_{1},\ldots,u_{n})):=\sum_{\mathbf{p}\in\tilde{S}_{\mathbf{f}}}{\rm ord}^{+}_{\mathbf{p}}(\widetilde{G}(u_{1},\ldots,u_{n}))\leq\frac{(n+1)\tilde{c}}{\ell}\max_{0\leq i\leq n}\{\deg f_{i}\}\leq\epsilon\deg\mathbf{f}$